# Sharp Thresholds for Temporal Motifs and Doubling Time in Random Temporal Graphs

### v1

@@ 8.1 Technical Preliminaries
The result for $R(P)$ holds trivially as $R(P)\geq 1$ and $|A|!\leq|A|^{|A|}$. For the other inequality, we will separate into two cases, and show that the minimum in each is at least constant.
Firstly, if $|A|>\beta$, since $|A|$ is finite, $\beta$ is integral and $\beta^{|A|}$ is positive, the infimum of $\inf_{\text{height(P)}\leq\beta<|A|}\frac{R(P,\beta)}{\beta^{|A|}}$ must be finite and non-negative. Furthermore, since $\beta>\text{height}(P)$, $R(P,\beta)>1$ as the assignment of each element to its height trivially satisfies $P$. Thus, if $|A|>\beta$,
@@
Secondly, we must consider the case where $|A|\leq\beta$. In order to obtain a bound, we will consider the strict total order $P^{\prime}$ produced by some arbitrary linear extension of $P$. It follows that $R(P,\beta)>R(P^{\prime},\beta)$ for all $\beta$, as any assignment satisfying $P^{\prime}$ must satisfy $P$. However, $R(P^{\prime},\beta)=\binom{\beta}{A}$ and thus for $|A|\leq\beta$,
@@
We will also need the following notion of automorphism for temporal motifs. In particular we require, that the mapping over the edges produced by the automorphism of the vertices corresponds to an automorphism of the associated partial order. The number of such automorphisms will play a crucial role in preventing over counting of the number of occurrences of temporal motifs when multiple occurrences share the same labels. See 4.2 Additionally, we require the following simple domination result, which relates the probability of a temporal motif appearing in two instances of our random graph models based on stochastic domination between their label multiplicity distributions. This will serve as an extremely useful tool which allows us to reduce the problem of obtaining certain bounds for arbitrary label multiplicity distributions to the case of Bernoulli or otherwise bounded label multiplicity distributions. See 4.3
@@
Thus, we have stochastic domination between the distribution of $A$ and $B$ and our claim.
@@
We define $\mathcal{S}$ to be the set of all occurrences of $H$ in the footprint graph $G=(V,E)$ and for any $S=(V_{S},E_{S})$ we define $\mathcal{T}_{S}$ to be the set of all mappings from $E_{S}\to\mathbb{N}_{0}$. Furthermore, we define the following order over $\cup_{I\subseteq G}\mathcal{T}_{I}$ where for $a\in\mathcal{T}_{S}$ and $b\in\mathcal{T}_{T}$ we describe $a\geq b$ if and only if $\forall e\in E_{S\cap T}$ $a(e)\geq b(e)$ and $S\cap T\neq\emptyset$. Define $\tau^{\text{max}}_{S}$ to be the random function such that $\forall e\in E_{S}$ we have $\tau^{\text{max}}_{S}(e)=l_{e}$ and $\tau^{\text{min}}_{S}$ to be the constant mapping $\tau^{\text{min}}_{S}(e)=1$. Additionally, for mathematical convenience we shall treat $\lambda(e)$ as a tuple ordered by the order of sampling rather than a multi-set and we shall assume that for each edge the value of all labels are sampled first, followed by how many will actually appear.
For a random sample $\mathcal{G}$ from $\Gamma_{[n]}(\psi)$, and $S\in\mathcal{S}$ and $\tau\in\mathcal{T}_{S}$ we define $A_{S,\tau}$ to be the indicator for the event $\{\forall e\in E_{S}:l_{e}\geq\tau(e)\}$ (i.e. the event that $\tau$ describes a set of labels that actually exist for $S$). Similarly, we define $B^{(\delta)}_{S,\tau}$ to be the indicator for the event that the labeling function $\lambda^{\prime}:E_{S}\to(0,1]$ given by $\lambda^{\prime}(e)=t_{e,\tau(e)}$ has the properties that $\max_{e_{1},e_{2}\in E_{S}}(\lambda^{\prime}(e_{1})-\lambda^{\prime}(e_{2}))\leq\delta$ and that $\lambda^{\prime}(e)$ is a valid realization of $P$.
@@
The indicator random variable $C^{(\delta)}_{S,\tau}=A_{S,\tau}B^{(\delta)}_{S,\tau}$ then corresponds to the indicator for the event that $(S,\tau)$ describes a real $\delta$-temporal $(H,P)$ motif of $\mathcal{G}$. Finally, $\mathbf{C}_{\delta}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}C^{(\delta)}_{S,\tau}$ corresponds to the total number of appearances of $(H,P)$ as a $\delta$-temporal motif of $\mathcal{G}$.
Analogously, for a random sample $\mathcal{H}$ from $\Gamma_{[n]}(\psi,T)$, which we obtain by taking the $T$-discretization of $\mathcal{G}$, we define $D^{(\beta),T}_{S,\tau}$ to be the indicator for the event that the labeling function $\lambda^{\prime}:E_{S}\to(0,1]$ given by $\lambda^{\prime}(e)=t_{e,\tau(e)}$ has the properties that $\max_{e_{1},e_{2}\in E_{S}}(\lambda^{\prime}(e_{1})-\lambda^{\prime}(e_{2}))<\delta$ and that $\lambda^{\prime}(e)$ is a valid realization of $P$. The indicator event $F^{(\beta),T}_{S,\tau}=A_{S,\tau}D^{(\beta),T}_{S,\tau}$ corresponds to the indicator for the event that $(S,\tau)$ describes a real $\beta$-temporal $(H,P)$ motif of $\mathcal{H}$. Finally, $\mathbf{F}_{(\beta),T}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}F^{(\beta),T}_{S,\tau}$ corresponds to the total number of appearances of $(H,P)$ as a $\beta$-temporal motif of $\mathcal{H}$.
@@
Now note that, whether or not a given set of labels fall within a specific interval, is independent of the order they fall in. Thus we immediately obtain the following more general statement.
@@
Unfortunately, the precise characterization of the discrete variant $D^{(\beta),T}_{S,\tau}$ is considerably less nice to work with. It is in fact significantly more convenient (and as we will later see, sufficient) to work with the following loose bounds derived from the continuous case. Consider that all labels that fall within an interval of length $\beta$ in $\Gamma_{[n]}(\psi,T)$ must fall within an interval of length $\frac{\beta}{T}$ in the associated instance of $\Gamma_{[n]}(\psi)$. Formally, we make the following observation. See 4.3 We can now apply this alongside, Lemma˜8.4, Corollary˜8.5 and Definition˜2.5 to obtain bounds on $D^{(\beta),T}_{S,\tau}$.
@@
Let $H=(V_{H},E_{H})$ a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges and $T\geq\beta\geq 1$. For any $S\in\mathcal{S},\tau\in\mathcal{T}_{S}$,
@@
Proof 8.7.
Let $X_{S,\tau}$ be the event that the labels of $\tau$ fall within an interval of length $\beta$ and let $Y_{S,\tau}$ be the event that the labels of $\tau$ obey $P$. Now consider a sample from $\Gamma_{[V]}(\psi)$, it follows from Definition˜2.5, Lemma˜8.4 and the fact that the set of intervals of the form $\frac{i}{T},\frac{i+\beta-1}{T}]$ is a subset of the continuous intervals of length $\frac{\beta}{T}$ that,
@@
$\mathbb{P}{[G_{S,\tau}|X_{S,\tau}]}=\frac{R(P,\beta)}{\beta^{|E_{H}|}}.$
@@
For the lower bound, we consider sampling $\mathcal{H}$ first from $\Gamma_{V}(\psi)$ and then taking the $T$-discretization of our sample to obtain our dependent sample $\mathcal{G}$ from $\Gamma_{V}(\psi,\tau)$ in accordance with Definition˜2.5. Now let $Z_{S,\tau}$ be the event that the labels of $\tau$ fall within an interval of length $\frac{\beta}{2T}$ in $\mathcal{H}$. From Lemma˜8.4, we have that
@@
Finally, for the case where $\delta=1$, there are $T$ possible snapshots and all labels land in each one with probability $T^{-|E_{H}|}$. Since the labels falling each snapshot are mutually exclusive, we conclude with a tight union bound.
@@ 8.1.1 First Moment
Proof 8.8.
We present the proof for the continuous case, and defer the proof of the discrete cases to the appendices, as they follow from near identical calculations. By independence we have that $\mathbb{P}{\left[A_{S,\tau}=1\right]}=\prod_{e\in E}\mathbb{P}{\left[\tau(e)\leq l_{e}\right]}$. By linearity, we obtain,
@@
From Corollary˜8.5 we have,
@@
Since $\sum_{\tau\in\mathcal{T}_{S}}\mathbb{P}{[\tau(e)\leq l_{e}]}=r$ is both finite and absolutely convergent, by the repeated application of distributivity over absolutely convergent sequences,
@@
With that established, we may turn our intention to construction of the actual bound. The observant reader may notice that the expectations obtained above will tend to zero whenever $\delta(n)=o(n^{-\rho_{H}})$ and $|V_{H}|$ is fixed. Infact Lemma˜4.4 is sufficient to show the lower bounds of Theorem˜3.1 and Theorem˜3.2 via a straightforward application of the first moment method. Unfortunately, the same cannot be said for Theorem˜3.4. Indeed, even a cursory inspection of the problem will reveal that $C^{(\delta)}_{S_{1},\tau_{1}}$ and $C^{(\delta)}_{S_{2},\tau_{2}}$ have a strong positive association whenever $\tau_{1}$ and $\tau_{2}$ share labels or even when $S_{1}$ and $S_{2}$ share a single edge. Consequently, the first moment produces an extremely weak bound that explodes whenever $r>1$ and the size of $H$ is unbounded. In order to control these dependencies and limit the contribution of additional occurrences after the first, we make use of a slightly more sophisticated bound based on the ratio of the expectation and the expectation conditioned on being non-zero. See 4.5
Proof 8.9.
@@
However, we still must determine $\hat{\tau}$. First, note that, for all $\tau_{1},\tau_{2}\in\mathcal{T}_{S}$, $\mathbb{P}{[B^{(\delta)}_{S,\tau_{1}}=1|B^{(\delta)}_{S,\tau_{2}}=1]}$ depends only on $\{e\in E_{S}:\tau_{1}(e)=\tau_{2}(e)\}$ the set of edge labels shared by the two. Now for any $\tau_{1},\tau_{2},\tau_{3}\in\mathcal{T}_{S}$ where $\tau_{1},\tau_{2}\leq\tau_{3}$ we have by symmetry in the number of pairs of labelings sharing a given set of edge labels.
@@
The last expectation is trivially lower bounded by the probability that all labels yet to be sampled fall into a specific window of length $\delta$ and that all labels fall in a permitted order. Therefore
@@
$\mathbb{P}{[\mathbf{C}_{\delta,S}>0]}=\frac{\mathbb{E}[\mathbf{C}_{\delta,S}]}{\mathbb{E}[\mathbf{C}_{\delta,S}|\mathbf{C}_{\delta,S}\geq 1]}\geq\frac{\delta^{|E_{H}|}r^{|E_{H}|}\left(1+\frac{|E_{H}|(1-\delta)}{\delta}\right)}{\left(\frac{(1-\delta)q+\delta r}{q}\right)^{|E_{H}|}}$
@@ 8.1.2 Second Moment
Once again we wish to avoid a direct calculation of a lower bound on the probability of a given motif occurring, and so our strategy will be to use the second moment method. Unfortunately, we do not find a unified over-arching bound sufficient to obtain our results for both fixed and growing motifs, as positive association rears its head once more (although this time as an obstacle for the fixed case). In particular, when two potential occurrences of a motif share a footprint they remain dependent even if they do not share any labels due to the inherent randomness of the label multiplicity distribution. This presents an issue whenever $\psi$ is a non-degenerate non-Bernoulli distribution, as we will retain terms in our summations corresponding to such pairs. These terms contribute an additional $\delta$ factor to our calculations, which can be absorbed in the case of Theorem˜3.4 where $\delta$ is constant but a naive application of the second moment would lead to asymptotically weaker bounds in the case of Theorem˜3.1 and Theorem˜3.2 where $\delta$ is decreasing. In order to avoid this pitfall, we shall make use of two bounds, one for the case where $\psi$ is a Bernoulli random variable and one for the more general but still bounded case. We shall warm up with the considerably easier former case. See 4.6
Proof 8.12.
We will present the proof only the proof for the continuous case here, however it should be noted that the discrete case is identical replacing only the continuous events for their discrete analogues and $\delta$ with $\frac{\beta}{T}$.
@@
Applying, Corollary˜8.5 we have that,
@@
For $k=|\{e\in E_{S_{2}}:\tau(e)=1\}|$,
$\mathbb{P}{[B^{(\delta)}_{S_{1},\tau^{\text{min}}_{S_{1}}}=1,B^{(\delta)}_{S_{2},\tau}=1]}\leq\delta^{2|E_{H}|-k}\left(1+\frac{|E_{H}|(1-\delta)}{\delta}\right)^{2}$
Proof 8.15.
From Corollary˜8.5,
@@
We consider a relaxation of the event $\{B^{(\delta)}_{S_{2},\tau}=1|B^{(\delta)}_{S_{1},\tau^{\text{min}}_{S_{1}}}=1\}$. Let $Y$ be the indicator random variable for the event that all remaining labels of $\tau$ fall within $\delta$ of each other. It follows from Lemma˜8.4, that $\mathbb{P}{[Y|B^{(\delta)}_{S_{1},\tau^{\text{min}}_{S_{1}}}=1]}\leq\delta^{|E_{H}|-k}\left(1+\frac{|E_{H}|(1-\delta)}{\delta}\right)$. Thus,

### v2

@@ 8.1 Technical Preliminaries
The result for $R(P)$ holds trivially as $R(P)\geq 1$ and $|A|!\leq|A|^{|A|}$. For the other inequality, we will separate into two cases, and show that the minimum in each is at least constant. Since $\beta>\text{height}(P)$, $R(P,\beta)>1$ (as the assignment of each element to its height trivially satisfies $P$). Thus, if $|A|>\beta$,
@@
Secondly, we must consider the case where $|A|\leq\beta$. In order to obtain a bound, we will consider the strict total order $P^{\prime}$ produced by some arbitrary linear extension of $P$. It follows that $R(P,\beta)>R(P^{\prime},\beta)$ for all $\beta$, as any assignment satisfying $P^{\prime}$ must satisfy $P$. However, $R(P^{\prime},\beta)=\binom{\beta}{|A|}$ and thus for $|A|\leq\beta$,
@@
We will also need the following notion of automorphism for temporal motifs. In particular, we require that the mapping over the edges produced by the automorphism of the vertices corresponds to an automorphism of the associated partial order. The number of such automorphisms will play a crucial role in preventing the over-counting of the number of occurrences of temporal motifs when multiple occurrences share the same labels. See 4.2 Additionally, we require the following simple domination result, which relates the probability of a temporal motif appearing in two instances of our random graph models based on stochastic domination between their label multiplicity distributions. This will serve as an extremely useful tool which allows us to reduce the problem of obtaining certain bounds for arbitrary label multiplicity distributions to the case of Bernoulli or otherwise bounded label multiplicity distributions. See 4.3
@@
Thus, we have a stochastic order between the distribution of $A$ and $B$ and our claim.
@@
We define $\mathcal{S}$ to be the set of all occurrences of $H$ in the footprint graph $G=(V,E)$ and for any $S=(V_{S},E_{S})$ we define $\mathcal{T}_{S}$ to be the set of all mappings from $E_{S}\to\mathbb{N}_{0}$. Furthermore, we define the following order over $\cup_{I\subseteq G}\mathcal{T}_{I}$ where for $a\in\mathcal{T}_{S}$ and $b\in\mathcal{T}_{T}$ we describe $a\geq b$ if and only if $\forall e\in E_{S\cap T}$ $a(e)\geq b(e)$ and $S\cap T\neq\emptyset$. Define $\tau^{\text{max}}_{S}$ to be the random function such that $\forall e\in E_{S}$ we have $\tau^{\text{max}}_{S}(e)=l_{e}$ and $\tau^{\text{min}}_{S}$ to be the constant mapping $\tau^{\text{min}}_{S}(e)=1$. Additionally, for mathematical convenience we shall treat $\lambda(e)$ as a tuple ordered by the order of sampling rather than a multi-set and we shall assume that for each edge the value of all labels are sampled first, followed by how many will actually appear. We will now define a number of useful random variables:
Definition 8.4.
For a random sample $\mathcal{G}$ from $\Gamma_{[n]}(\psi)$, and $S\in\mathcal{S}$ and $\tau\in\mathcal{T}_{S}$ we define:
  • $A_{S,\tau}$ to be the indicator for the event $\{\forall e\in E_{S}:l_{e}\geq\tau(e)\}$ (i.e. the event that $\tau$ describes a set of labels that actually exist for $S$)
  • $B^{(\delta)}_{S,\tau}$ to be the indicator for the event that the labeling function $\lambda^{\prime}:E_{S}\to(0,1]$, mapping $e$ to its $\tau(e)$th sampled label has both of the following properties:
    1. 1.
      $\max_{e_{1},e_{2}\in E_{S}}(\lambda^{\prime}(e_{1})-\lambda^{\prime}(e_{2}))\leq\delta$ (i.e. all labels occur within a window of $\delta$)
    2. 2.
      $\lambda^{\prime}(e)$ is a valid realization of $P$.
@@
  • $C^{(\delta)}_{S,\tau}=A_{S,\tau}B^{(\delta)}_{S,\tau}$ (i.e. the indicator for the event that $(S,\tau)$ describes a genuine $\delta$-temporal $(H,P)$ motif in $\mathcal{G}$).
  • $\mathbf{C}_{\delta}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}C^{(\delta)}_{S,\tau}$ (i.e. the total number of appearances of $(H,P)$ as a $\delta$-temporal motif of $\mathcal{G}$)
Analogously, for the discrete case,
Definition 8.5.
For a random sample $\mathcal{H}$ from $\Gamma_{[n]}(\psi,T)$ (obtained by taking the $T$ discretization of $\mathcal{G}$), and $S\in\mathcal{S}$ and $\tau\in\mathcal{T}_{S}$ we define:
  • $D^{(\beta),T}_{S,\tau}$ to be the indicator for the event that the labeling function $\lambda^{\prime}:E_{S}\to(0,1]$ mapping $e$ to its $\tau(e)$th label has both of the following properties:
    1. 1.
      $\max_{e_{1},e_{2}\in E_{S}}(\lambda^{\prime}(e_{1})-\lambda^{\prime}(e_{2}))<\beta$ (i.e. all labels fall within a window of $\beta$)
    2. 2.
      $\lambda^{\prime}(e)$ is a valid realization of $P$
  • $F^{(\beta),T}_{S,\tau}=A_{S,\tau}D^{(\beta),T}_{S,\tau}$ (i.e. the indicator for the event that $(S,\tau)$ describes a genuine $\beta$-temporal motif of $\mathcal{H}$).
  • $\mathbf{F}_{(\beta),T}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}F^{(\beta),T}_{S,\tau}$ (i.e. the total number of appearances of $(H,P)$ as a $\beta$ temporal motif in $\mathcal{H}$).
@@
Now note that, whether or not a given set of labels all fall within a specific interval, is independent of the order they fall in. Thus we immediately obtain the following more general statement.
@@
Unfortunately, the precise characterization of the discrete variant $D^{(\beta),T}_{S,\tau}$ is considerably less nice to work with. It is in fact significantly more convenient (and as we will later see, sufficient) to work with the following loose bounds derived from the continuous case. Consider that all labels that fall within an interval of length $\beta$ in $\Gamma_{[n]}(\psi,T)$ must fall within an interval of length $\frac{\beta}{T}$ in the associated instance of $\Gamma_{[n]}(\psi)$. Formally, we make the following observation. See 4.3 We can now apply this alongside, Lemma˜8.6, Corollary˜8.7 and Definition˜2.5 to obtain bounds on $D^{(\beta),T}_{S,\tau}$.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges and $T\geq\beta\geq 1$. For any $S\in\mathcal{S},\tau\in\mathcal{T}_{S}$,
@@
Proof 8.9.
Let $X_{S,\tau}$ be the event that the labels of $\tau$ fall within an interval of length $\beta$ and let $Y_{S,\tau}$ be the event that the labels of $\tau$ obey $P$. Now consider a sample from $\Gamma_{[V]}(\psi)$, it follows from Definition˜2.5, Lemma˜8.6 and the fact that the set of intervals of the form $(\frac{i}{T},\frac{i+\beta}{T}]$ is a subset of the continuous intervals of length $\frac{\beta}{T}$ that,
@@
$\mathbb{P}{[Y_{S,\tau}|X_{S,\tau}]}=\frac{R(P,\beta)}{\beta^{|E_{H}|}}.$
@@
For the lower bound, we consider sampling $\mathcal{H}$ first from $\Gamma_{V}(\psi)$ and then taking the $T$-discretization of our sample to obtain our dependent sample $\mathcal{G}$ from $\Gamma_{V}(\psi,\tau)$ in accordance with Definition˜2.5. Now let $Z_{S,\tau}$ be the event that the labels of $\tau$ fall within an interval of length $\frac{\beta}{2T}$ in $\mathcal{H}$. From Lemma˜8.6, we have that
@@
Finally, for the case where $\delta=1$, there are $T$ possible snapshots and all labels land in each one with probability $T^{-|E_{H}|}$. Since the labels falling in each snapshot are mutually exclusive, we conclude with a tight union bound.
@@ 8.1.1 First Moment
Proof 8.10.
We present the proof for the continuous case, and defer the proof of the discrete cases to the appendices, as they follow from near identical calculations. By linearity, we obtain,
@@
From Corollary˜8.7 we have,
@@
Since $\sum_{\tau\in\mathcal{T}_{S}}\mathbb{P}{[\tau(e)\leq l_{e}]}=r$ is both finite and absolutely convergent, we observe that this is equivalent to the following via the $|E_{H}|$ fold Cauchy product,
@@
With that established, we may turn our attention to construction of the actual bound. The observant reader may notice that the expectations obtained above will tend to zero whenever $\delta(n)=o(n^{-\rho_{H}})$ and $|V_{H}|$ is fixed. Infact Lemma˜4.4 is sufficient to show the lower bounds of Theorem˜3.1 and Theorem˜3.2 via a straightforward application of the first moment method. Unfortunately, the same cannot be said for Theorem˜3.4. Indeed, even a cursory inspection of the problem will reveal that $C^{(\delta)}_{S_{1},\tau_{1}}$ and $C^{(\delta)}_{S_{2},\tau_{2}}$ have a strong positive association whenever $\tau_{1}$ and $\tau_{2}$ share labels or even when $S_{1}$ and $S_{2}$ share a single edge. Consequently, the first moment produces an extremely weak bound that explodes whenever $r>1$ and the size of $H$ is unbounded. In order to control these dependencies and limit the contribution of additional occurrences after the first, we make use of a slightly more sophisticated bound based on the ratio of the expectation and the expectation conditioned on being non-zero. See 4.5
Proof 8.11.
@@
However, we still must determine $\hat{\tau}$. First, note that, for all $\tau_{1},\tau_{2}\in\mathcal{T}_{S}$, $\mathbb{P}{[B^{(\delta)}_{S,\tau_{1}}=1|B^{(\delta)}_{S,\tau_{2}}=1]}$ depends only on $\{e\in E_{S}:\tau_{1}(e)=\tau_{2}(e)\}$ (the set of edge labels shared by the two). Now for any $\tau_{1},\tau_{2},\tau_{3}\in\mathcal{T}_{S}$ where $\tau_{1},\tau_{2}\leq\tau_{3}$ we have, by the symmetry in the number of pairs of labelings sharing a given set of edge labels, that,
@@
The last expectation is trivially lower bounded by the probability that all labels yet to be sampled fall into a specific window of length $\delta$ and that all labels fall in a permitted order. Therefore,
@@
$\mathbb{P}{[\mathbf{C}_{\delta,S}>0]}=\frac{\mathbb{E}[\mathbf{C}_{\delta,S}]}{\mathbb{E}[\mathbf{C}_{\delta,S}|\mathbf{C}_{\delta,S}\geq 1]}\leq\frac{\delta^{|E_{H}|}r^{|E_{H}|}\left(1+\frac{|E_{H}|(1-\delta)}{\delta}\right)}{\left(\frac{(1-\delta)q+\delta r}{q}\right)^{|E_{H}|}}$
@@ 8.1.2 Second Moment
Once again we wish to avoid a direct calculation of a lower bound on the probability of a given motif occurring, and so our strategy will be to use the second moment method. Unfortunately, we do not find a unified overarching bound sufficient to obtain our results for both fixed and growing motifs, as positive association rears its head once more (although this time as an obstacle for the fixed case). In particular, when two potential occurrences of a motif share a footprint they remain dependent even if they do not share any labels, due to the inherent randomness of the label multiplicity distribution. This presents an issue whenever $\psi$ is a non-degenerate non-Bernoulli distribution, as we will retain terms in our summations corresponding to such pairs. These terms contribute an additional $\delta$ factor to our calculations, which can be absorbed in the case of Theorem˜3.4 where $\delta$ is constant but a naive application of the second moment would lead to asymptotically weaker bounds in the case of Theorem˜3.1 and Theorem˜3.2 where $\delta$ is decreasing. In order to avoid this pitfall, we shall make use of two bounds, one for the case where $\psi$ is a Bernoulli random variable and one for the more general but still bounded case. We shall warm up with the considerably easier former case. See 4.6
Proof 8.14.
We will present only the proof for the continuous case here, however it should be noted that the discrete case is identical replacing only the continuous events for their discrete analogues and $\delta$ with $\frac{\beta}{T}$.
@@
Applying, Corollary˜8.7 we have that,
@@
For $k=|\{e\in E_{S_{2}}:\tau(e)\neq 1\}|$,
Proof 8.17.
From Corollary˜8.7,
@@
We consider a relaxation of the event $\{B^{(\delta)}_{S_{2},\tau}=1|B^{(\delta)}_{S_{1},\tau^{\text{min}}_{S_{1}}}=1\}$. Let $Y$ be the indicator random variable for the event that all remaining labels of $\tau$ fall within $\delta$ of each other. It follows from Lemma˜8.6, that $\mathbb{P}{[Y|B^{(\delta)}_{S_{1},\tau^{\text{min}}_{S_{1}}}=1]}\leq\delta^{k}\left(1+\frac{|E_{H}|(1-\delta)}{\delta}\right)$. Thus,